OpenStudy (anonymous):

help please (cosx-sqrt2/2)(secx-1)=0 use reciprocal identity to express the equation involving secant in term of sine,cosine,or tangent.

OpenStudy (anonymous):

\[\frac{1}{\cos \theta}=\sec \theta\]

OpenStudy (anonymous):

so thats the answer??

OpenStudy (anonymous):

No. That's what you use to get the answer.

OpenStudy (anonymous):

so 1/cosx=sec?

OpenStudy (anonymous):

Just substitute \(\frac{1}{\cos \theta}\) for \(\sec \theta\) in your equation and solve.

OpenStudy (anonymous):

sqrt2/sec?

OpenStudy (anonymous):

-1/2(sectheta-1)(sqrt2-2cosx=0

OpenStudy (anonymous):

@Limitless is that correct?

OpenStudy (campbell_st):

is is quite simple its a factorised equation so ou need to solve \[\cos(x) - \frac{\sqrt{2}}{2}= 0 \] \[\cos(x) = \frac{\sqrt{2}}{2}\] \[x = {\pi}{4} ..or... \frac{7\pi}{4} ..... or x = 45 ..or... 315\] as well as \[\sec(x) - 1 = 0\] \[\frac{1}{\cos(x)} = 1 ... .... \cos(x) = 1 .... then .... x = 0, 2\pi ...or x = 0, 360\]

OpenStudy (anonymous):

Sorry, I was AFK. One moment.

OpenStudy (anonymous):

\[ \begin{align} \left(\cos (x)-\sqrt{\frac{2}{2}}\right)\left(\sec (x)-1\right)&=\cos(x)\left(\sec (x)-1\right)-\sqrt{\frac{2}{2}}\left(\sec (x)-1\right)\\ &=\cos(x)\sec(x)-\cos(x)-\sqrt{\frac{2}{2}}\sec(x)+\sqrt{\frac{2}{2}}\\ &=\cos(x)\frac{1}{\cos(x)}-\cos(x)-\sqrt{\frac{2}{2}}\frac{1}{\cos(x)}+\sqrt{\frac{2}{2}}\\ &=1-\cos(x)-\sqrt{\frac{2}{2}}\frac{1}{\cos(x)}+\sqrt{\frac{2}{2}}\\ &=-\cos(x)-\frac{\sqrt{\frac{2}{2}}}{\cos(x)}+1+\sqrt{\frac{2}{2}}\\ \text{ or, if you prefer, } &=-\cos(x)-\frac{\sqrt{4}}{2\cos(x)}+\frac{2+\sqrt{4}}{2} \end{align} \]